OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = x + 5 and that pass through the point (-2, -1).

OpenStudy (anonymous):

A)parallel: y = -x - 1 perpendicular: y = x + 2 B)parallel: y = x - 1 perpendicular: y = -x + 1 C)parallel: y = x + 1 perpendicular: y = -x - 3 D)parallel: y = 2x - 2 perpendicular: y = -2x - 1